# Direct Neural Machine Translation with Task-level Mixture of Experts models

### v1

@@ Abstract
Direct neural machine translation (direct NMT) is a type of NMT system that translates text between two non-English languages. Direct NMT systems often face limitations due to the scarcity of parallel data between non-English language pairs. Several approaches have been proposed to address this limitation, such as multilingual NMT and pivot NMT (translation between two languages via English). Task-level Mixture of expert models (Task-level MoE), an inference-efficient variation of Transformer-based models, has shown promising NMT performance for a large number of language pairs. In Task-level MoE, different language groups can use different routing strategies to optimize cross-lingual learning and inference speed. In this work, we examine Task-level MoE’s applicability in direct NMT and propose a series of high-performing training and evaluation configurations, through which Task-level MoE-based direct NMT systems outperform bilingual and pivot-based models for a large number of low and high-resource direct pairs, and translation directions. Our Task-level MoE with 16 experts outperforms bilingual NMT, Pivot NMT models for 7 language pairs, while pivot-based models still performed better in 9 pairs and directions.
@@ 1 Introduction
Figure 1: Task-level MoE model, with LP - based routing; each Language Pair is routed through a top-2 router to an expert in the model experts’ layer. From a pretrained Task-level MoE model we can extract a smaller dense network specializing in a certain task, e.g. ja-ko Neural Machine Translation.
Neural machine translation (NMT) (Sutskever et al., 2014; Bahdanau et al., 2014; Firat et al., 2016a; Johnson et al., 2017; Zoph and Knight, 2016; Bahdanau et al., 2014) has made remarkable progress in recent years, particularly with the advent of the Transformer architecture (Vaswani et al., 2017), which has shown impressive results in bilingual and multilingual translation tasks (Aharoni et al., 2019; Hassan et al., 2018; Tang et al., 2021). Large Multilingual Neural Machine Translation models in particular have shown noteworthy performance and can serve translations between any languages, even for language pairs not seen in training, for which though NMT quality is often still low. The focus so far has primarily been on English-centric NMT, as there is an abundance of English-centric MT data (Tiedemann, 2018) but few parallel data for non-English (low-resource) language pairs. Hence the need for improving methods for Direct NMT, i.e. NMT between non-English languages has become evident (Yang et al., 2021; Zhang et al., 2020).
@@
Sparse models, and specifically, sparsely activated Mixture-of-Experts (MoE) models (Shazeer et al., 2017; Lepikhin et al., 2020; Zoph et al., 2022; Ryabinin and Gusev, 2020), shape a promising direction in creating a universal translation model between all languages, while addressing efficiency concerns. We are the first to investigate performance of a specific type of MoE models, Task-level MoE models (kudugunta-etal-2021-beyond-distillation) in the context of Direct Neural Machine Translation (NMT), focusing on the impact of different task-to-expert mapping methods on model performance and expert utilization. In our experiments, we train large multilingual Task-level MoE models across various configurations, including Zero-shot, Zero-resource, and scenarios where direct language pair data is available. We conduct a comparative analysis of BLEU scores across Task-level MoE models with different mapping strategies and expert counts, as well as bilingual NMT and pivot-based NMT models; our findings serve as a valuable resource for determining the optimal model configuration to achieve high-quality translation results in various direct language pairs. By training large sparse multilingual Task-level MoE models which perform well for direct pairs, and showcase training efficiency, we may be able to extract smaller expert-specific dense models, which specialize in translating between Language Pairs or translating into one particular language. These models are ready for deployment and can replace bilingual or pivot-based models (Figure 1). Prior work on Task-level MoEs kudugunta-etal-2021-beyond-distillation focused on demonstrating the desired inference efficiency of task-level routing, compared to other routing approaches (token- or sentence-level), for the task of Multilingual Neural Machine Translation (MNMT) between English-centric pairs (En-yy and xx–En). To the best of our knowledge, we’re the first ones to examine performance of Task-level MoE models in Direct NMT.
@@
2 Related Work
Direct Neural Machine Translation
Firat et al. (2016b) first proposed a multilingual multi-way zero-resource NMT model which outperformed previous pivot-based approaches, laying the foundation for numerous subsequent zero-shot and/or zero-resource NMT methods. Various researchers have since expanded upon this foundational work, exploring different techniques and strategies for improving direct NMT (Ji et al., 2020; Johnson et al., 2017; Zhang et al., 2020; Chen et al., 2017; Lu et al., 2018; Al-Shedivat and Parikh, 2019; Rios et al., 2020; Zheng et al., 2017; Gu et al., 2019; Arivazhagan et al., 2019; Kim et al., 2019; Cheng and Cheng, 2019; Chen et al., 2018; Currey and Heafield, 2019; Ballesteros and Sanderson, 2003; Lakew et al., 2021). More specifically, Freitag and Firat (2020) constructed a multi-way aligned training set from existing training corpora, aligning examples in different language pairs with identical source or target sentences. They introduced a hierarchical data sampling strategy to prevent the over-representation of English in the training set and employed a temperature-based strategy to select target languages, while choosing source languages uniformly. Their complete Multilingual Neural Machine Translation models demonstrated improved performance on direct pairs and comparable to baseline results on English-centric pairs. Liu et al. (2020) discretized the encoder output latent space into entries in a learned codebook, enabling the representation of source sentences in a common language and increasing model robustness and performance in zero-shot setups. More recently, eriguchi-etal-2022-building pre-trained and fine-tuned a many-to-many multilingual NMT model that surpassed several pivoting and bilingual baselines for Direct NMT, and ElNokrashy et al. (2022), presented a novel approach in which source and target language tokens are appended to the encoder input, while target language tokens are added to the decoder input. This differs from previous methods that used Source-only tokens and beginning of Sentence tokens for the encoder and decoder inputs, respectively, and aims to improve performance of Direct NMT models by providing more relevant context to both the encoder and decoder.
bg–mk fr–ar ja–th zh–ko ja–vi ja–ko ja–zh $\rightarrow$ $\leftarrow$ $\rightarrow$ $\leftarrow$ $\rightarrow$ $\leftarrow$ $\rightarrow$ $\leftarrow$ $\leftarrow$ $\rightarrow$ $\leftarrow$ Task-level MoE (LP) lp_a 19.18 - - 16.33 17.50 21.34 19.08 19.18 25.3 25.25 22.52 lp_b 19.07 4.47 2.13 0.44 2.79 2.05 2.51 0.47 0.96 4.46 0.71 lp_c 19.04 2.69 1.18 0.36 0.50 0.59 0.47 0.45 0.80 2.05 0.56 Task-level MoE (TL) tl_a 20.29 7.37 5.38 8.30 25.33 12.61 20.27 5.95 14.25 26.39 0.01 tl_b 20.35 7.32 5.50 8.33 25.36 12.43 20.34 5.96 14.26 26.40 10.49 Bilingual 21.28 9.71 19.60 6.24 7.26 5.71 3.16 8.20 6.12 4.47 3.16 Pivot-level 27.03 14.91 21.57 18.34 39.28 28.13 35.24 18.97 21.49 36.69 19.98 Number of train sentences 205,651 17,163,359 6,200 498,968 604,940 974,896 1,339,622
@@ 5 Conclusions
We conducted a thorough analysis of the usage of Task-level Mixture of Expert models for Direct Neural Machine Translation. Our experiments reveal the strengths and weaknesses of different approaches and shed light on which configurations work best for specific direct language pairs. Our comparisons help in identifying best-performing models and offer valuable insights into how varying numbers of experts, and different task-to-expert mapping methods, during training and inference, can influence direct pairs’ translation quality in Task-level MoE models. Specifically, we noticed that Task-level MoE NMT models, along with pivot-based approaches, are frequently top performers for numerous direct language pairs. However, their performance varies between different pairs and translation directions. Besides observing NMT quality results, our analysis of expert’s usage throughout training also serves as an informative way to visualize and understand mapping of language pairs to model experts in the encoder and decoder throughout training.

### v2

@@ Abstract
Direct neural machine translation (direct NMT) is a type of NMT system that translates text between two non-English languages. Direct NMT systems often face limitations due to the scarcity of parallel data between non-English language pairs. Several approaches have been proposed to address this limitation, such as multilingual NMT and pivot NMT (translation between two languages via English). Task-level Mixture of expert models (Task-level MoE), an inference-efficient variation of Transformer-based models, has shown promising NMT performance for a large number of language pairs. In Task-level MoE, different language groups can use different routing strategies to optimize cross-lingual learning and inference speed. In this work, we examine Task-level MoE’s applicability in direct NMT and propose a series of high-performing training and evaluation configurations, through which Task-level MoE-based direct NMT systems outperform bilingual and pivot-based models for a large number of low and high-resource direct pairs, and translation directions. Our Task-level MoE model with 16 experts outperforms bilingual NMT, Pivot NMT models for 7 language pairs, while pivot-based models still performed better in 9 pairs and directions.
@@ 1 Introduction
Figure 1: Task-level MoE model, with LP - based routing; each Language Pair is routed through a top-2 router to an expert in the model experts’ layer. From a pretrained Task-level MoE model we can extract a smaller dense network specializing in a certain task, e.g. ja-ko NMT.
NMT (Sutskever et al., 2014; Bahdanau et al., 2014; Firat et al., 2016a; Johnson et al., 2017; Zoph and Knight, 2016; Bahdanau et al., 2014) has made remarkable progress in recent years, particularly with the advent of the Transformer architecture (Vaswani et al., 2017), which has shown impressive results in bilingual and multilingual translation tasks (Aharoni et al., 2019; Hassan et al., 2018; Tang et al., 2021). Large Multilingual NMT models in particular have shown noteworthy performance and can serve translations between any languages, even for language pairs not seen in training, for which though NMT quality is often still low. The focus so far has primarily been on English-centric NMT, as there is an abundance of English-centric MT data (Tiedemann, 2018) but few parallel data for non-English (low-resource) language pairs. Hence the need for improving methods for Direct NMT, i.e. NMT between non-English languages is evident (Yang et al., 2021; Zhang et al., 2020).
@@
Sparse models, and specifically, sparsely activated Mixture-of-Experts (MoE) models (Shazeer et al., 2017; Lepikhin et al., 2020; Zoph et al., 2022; Ryabinin and Gusev, 2020), shape a promising direction in creating a universal translation model between all languages, while addressing efficiency concerns. We are the first to investigate performance of a specific type of MoE models, Task-level MoE models (Kudugunta et al., 2021) in the context of Direct NMT, focusing on the impact of different task-to-expert mapping methods on model performance and expert utilization. In our experiments, we train large multilingual Task-level MoE models across various configurations, including Zero-shot, Zero-resource, and scenarios where direct language pair data is available. We conduct a comparative analysis of BLEU scores across Task-level MoE models with different mapping strategies and expert counts, as well as bilingual NMT and pivot-based NMT models; our findings serve as a valuable resource for determining the optimal model configuration to achieve high-quality translation results in various direct language pairs. By training large sparse multilingual Task-level MoE models which perform well for direct pairs, and showcase training efficiency, we may be able to extract smaller expert-specific dense models, which specialize in translating between Language Pairs or translating into one particular language. These models are ready for deployment and can replace bilingual or pivot-based models (Figure 1). Prior work on Task-level MoEs Kudugunta et al. (2021) demonstrates the inference efficiency of task-level routing, compared to other approaches (token- or sentence-level), for the task of Multilingual NMT between English-centric pairs (En-yy and xx–En). To the best of our knowledge, we’re the first ones to examine performance of Task-level MoE models in Direct NMT, offering a practical solution for expanding the coverage and quality of NMT across diverse language pairs.
2 Related Work
Direct Neural Machine Translation
Firat et al. (2016b) first proposed a multilingual multi-way zero-resource NMT model which outperformed previous pivot-based approaches, laying the foundation for numerous subsequent zero-shot and/or zero-resource NMT methods. Various researchers have since expanded upon this foundational work, exploring different techniques and strategies for improving direct NMT (Ji et al., 2020; Johnson et al., 2017; Zhang et al., 2020; Chen et al., 2017; Lu et al., 2018; Al-Shedivat and Parikh, 2019; Rios et al., 2020; Zheng et al., 2017; Gu et al., 2019; Arivazhagan et al., 2019; Kim et al., 2019; Cheng and Cheng, 2019; Chen et al., 2018; Currey and Heafield, 2019; Ballesteros and Sanderson, 2003; Lakew et al., 2021). More specifically, Freitag and Firat (2020) constructed a multi-way aligned training set from existing training corpora, aligning examples in different language pairs with identical source or target sentences. They introduced a hierarchical data sampling strategy to prevent the over-representation of English in the training set and employed a temperature-based strategy to select target languages, while choosing source languages uniformly. Their complete Multilingual NMT models demonstrated improved performance on direct pairs and comparable to baseline results on English-centric pairs. Liu et al. (2020) discretized the encoder output latent space into entries in a learned codebook, enabling the representation of source sentences in a common language and increasing model robustness and performance in zero-shot setups. More recently, eriguchi2022building pre-trained and fine-tuned a many-to-many multilingual NMT model that surpassed several pivoting and bilingual baselines for Direct NMT, and ElNokrashy et al. (2022), presented a novel approach in which source and target language tokens are appended to the encoder input, while target language tokens are added to the decoder input. This differs from previous methods that used Source-only tokens and beginning of Sentence tokens for the encoder and decoder inputs, respectively, and aims to improve performance of Direct NMT models by providing more relevant context to both the encoder and decoder.
@@ 2 Related Work
  ja-th bg-mk   fr-ar   zh–ko ja–vi   ja–ko ja–zh $\leftarrow$ $\rightarrow$ $\rightarrow$ $\leftarrow$ $\rightarrow$ $\leftarrow$ $\rightarrow$ $\leftarrow$ $\leftarrow$ $\rightarrow$ $\leftarrow$ Task-level MoE (LP) lp_a 16.33 17.50 19.18 - - 21.34 19.08 19.18 25.3 25.25 22.52 lp_b 0.44 2.79 19.07 4.47 2.13 2.05 2.51 0.47 0.96 4.46 0.71 lp_c 0.36 0.50 19.04 2.69 1.18 0.59 0.47 0.45 0.80 2.05 0.56 Task-level MoE (TL) tl_a 8.30 25.33 20.29 7.37 5.38 12.61 20.27 5.95 14.25 26.39 0.01 tl_b 8.33 25.36 20.35 7.32 5.50 12.43 20.34 5.96 14.26 26.40 10.49 Bilingual 6.24 7.26 21.28 9.71 19.60 5.71 3.16 8.20 6.12 4.47 3.16 Pivot-level 18.34 39.28 27.03 14.91 21.57 28.13 35.24 18.97 21.49 36.69 19.98 Number of train sentences 6,200 205,651 17,163,359 498,968 604,940 974,896 1,339,622
@@ 5 Conclusions
We conducted a thorough analysis of the usage of Task-level Mixture of Expert models for Direct NMT. Our experiments reveal the strengths and weaknesses of different approaches and shed light on which configurations work best for specific direct language pairs. Our comparisons help in identifying best-performing models and offer valuable insights into how varying numbers of experts, and different task-to-expert mapping methods, during training and inference, can influence direct pairs’ translation quality in Task-level MoE models. Specifically, we noticed that Task-level MoE NMT models, along with pivot-based approaches, are frequently top performers for numerous direct language pairs. However, their performance varies between different pairs and translation directions. Besides observing NMT quality results, our analysis of expert’s usage throughout training also serves as an informative way to visualize and understand mapping of language pairs to model experts in the encoder and decoder throughout training. Future work can focus on enhancing this model for broader language coverage, including other low-resource languages, to further improve translation quality and efficiency.